# How Post-Training Reshapes LLMs: A Mechanistic View on Knowledge, Truthfulness, Refusal, and Confidence

### v2

@@ 7 Confidence
Confidence of LLMs is represented by the probability associated with the decoded token. Post-trained models are known to have different confidence compared to base models (OpenAI, 2023), which is also revealed in their drastically different outputs to the same prompts. Understanding and calibrating model confidence is an important research direction. Recently, entropy neurons have been shown to be a mechanism of modulating confidence that is persistent across models (Gurnee et al., 2024; Stolfo et al., 2024). Entropy neurons help calibrate the model’s confidence. They have relatively high weight norms and a low composition with the model’s unembedding matrix, so they influence the model’s output logits without affecting the token ranking and which token will be predicted, working similarly to the temperature parameter. We aim to study whether the difference in confidence between base and post models is caused by the difference in entropy neurons.
Entropy neurons are identified by checking the weight norm and logit attribution. First, we compute the logit attribution for each neuron in the final MLP layer by projecting its output weights onto the vocabulary space through the unembedding matrix. This projection (Equation 4) approximates the neuron’s direct effect on the final prediction logits:
@@
where $W_{out}$ is the weight vector of the last MLP layer, $W_{U}$ is the unembedding matrix, and $\|\cdot\|_{\text{dim}=1}$ denotes a row-wise norm. We then calculate the variance of this normalized projection (LogitVar), where a low LogitVar value indicates a relatively balanced contribution across all vocabulary tokens rather than promoting specific tokens. Entropy neurons typically have both a large weight norm (ensuring they are influential) and a low LogitVar (indicating balanced contribution across vocabulary tokens). Our identification process first selects the top 25% of neurons with the largest weight norms, and from this subset, we identify the 10 neurons with the lowest LogitVar values as entropy neurons from the final MLP layer. This methodology follows established practices from prior work and captures neurons that modulate output entropy without significantly affecting token ranking.
In our analysis comparing BASE and POST models, we found substantial overlap in identified entropy neurons, with highly similar ratios of $\left|\frac{\text{weight norm}}{\log(\text{LogitVar})}\right|$ between models. We show the detailed results in Appendix E. These finding suggests that the confidence regulation mechanism of entropy neurons remains largely unchanged during post-training, indicating that the observed confidence calibration differences between BASE and POST models likely stem from more subtle mechanistic changes that require sophisticated interpretability tools beyond current entropy neuron analysis to fully understand.
@@ 8 Discussion and Conclusion
To achieve effective post-training, it is important to understand how it shapes LLMs internally. In this paper, we analyze its effect on LLM’s internal mechanisms from four representative perspectives. We discover that post-training does not alter knowledge-storage locations and truthfulness directions significantly, and adapts original knowledge representations while developing some new ones. However, post-training changes the refusal direction. We also find that the confidence difference brought by post-training cannot be attributed to entropy neurons, further works need to be done.
Our findings could also benefit many real-world applications. As we have shown, general abilities such as factual knowledge and the internal belief of truthfulness are mostly developed during pre-training and remain unchanged in post-training. Although post-training develops new knowledge representations, the forward transfer remains valid. For fixing mistakes or outdated knowledge, this allows us to conveniently and effectively transfer knowledge editing developed on a base model to its post model. We can also transfer the hidden probe of truthfulness learned from base or post to each other, benefiting model steering. In contrast, some internal mechanisms are significantly modified by post-training, such as refusing harmful instructions. In these areas, a valuable application is to transfer the newly acquired capabilities from the post model to the base model, making it efficient for the base model to obtain such ability (Kissane et al., 2024a).
Although we concentrated on four key perspectives, future work could extend our framework to more complex capabilities, such as reasoning and instruction-following. These areas present significant methodological challenges for existing interpretability tools. We also find that properly defining the instruction-following ability is tricky, and a suitable technique to interpret this ability and verify it on base is also non-trivial. Also, future work could utilize the analysis to improve the post-training effectiveness and efficiency.
@@ Appendix A Details on Datasets
The Tulu_extracted dataset is an in-distribution dataset for the Llama-3.1-8B sft and Mistral-7B-v0.3 sft models. In order to construct it, we use GPT-4o to extract 100 factual knowledge statements from the Tulu-SFT dataset that was used to fine-tune the sft models (Lambert et al., 2024). Then we use GPT-4o to generate a false statement for each true factual statement by changing the subject, object, or subject-object relation.
@@ B.2 Few-shot prompting
For each dataset, we select 2 true examples and 2 false examples to conduct four-shot prompting. We randomly select them from the dataset once, and then we fix them. The selected examples are shown in Table 7. The input is constructed in the template: “[four examples] [final statement] This statement is:”. To eliminate the influence of example order, we randomly perturb the four examples for every (true, false) statement pairs, so different pairs might have different example orders, but the true and false statements in a pair have the same example order. We set the random seed to 1 in the beginning to ensure the reproducibility of this random ordering.
@@ B.4 Supplementary Quantitative Results
Our main experiments follow the setting of Marks & Tegmark (2024). We ask the LLM to classify the truthfulness of a statement. This setup differs from the traditional causal tracing setup (Meng et al., 2022), which uses LLM to output the object corresponding to a given subject. We choose this setting because of the following considerations. First, this setting (e.g., ”The city of Toronto is in Canada. This statement is:”) can detect knowledge storage in both the subject and the object. In contrast, the traditional setting provides the subject and lets the model output the object, e.g., ”The city of Toronto is in”. It can only detect knowledge storage in the subject. Second, our setting can test a wider range of factual knowledge. The traditional setting evaluates the patching’s influence by examining the output logit of the correct object, so it must have a fixed correct answer, such as the country of a city. But in many datasets, such as larger_than, statements like ”86 is larger than 57” don’t have a fixed correct answer. Any number less than the subject is correct here.
@@ C.3 Probe Transfer Accuracy on Mistral Family
Table 12: Probe transfer accuracy ($\uparrow$) of Mistral-7B-v0.3 base and instruct tested on 6 truthfulness datasets. For each row, the datasets from the other 5 rows are used for training. $p_{model_{1}}\rightarrow h_{model_{2}}$ means using the probe trained on $model_{1}$ to classify truthfulness direction in $model_{2}$. Probe transfer shows little difference ($\Delta$) compared to the same-model probe.
@@ C.4 Probe Intervention Coefficient Choice
To assess the robustness of our findings to the choice of scaling factor, we extended our experiments beyond the default scalar setting ($\lambda=\pm 1$) used in Marks & Tegmark (2024). Prior work has shown that scaling can impact intervention effectiveness (Li et al., 2024), motivating a broader evaluation.
@@ C.5 Probe Intervention on Mistral Family
Table 14: Intervention effect ($\uparrow$) of intervention on Mistral-7B-v0.3 base and instruct tested on 6 truthful datasets. For each row, the datasets from the other 5 rows are used for training. $t_{model_{1}}\mapsto h_{model_{2}}$ means using the truthfulness direction in $model_{1}$ to intervene $model_{2}$. Transfer truthful interventions show small differences ($\Delta$).
@@ D.1 Refusal Scores
Following Arditi et al. (2024); Kissane et al. (2024a) We use refusal keywords to determine if a model refuses an input. The keywords are listed in Figure 5. Refusal score on a dataset is calculated as the percentage of outputs of the model that contains these keywords at the beginning.
@@ D.3 Refusal Direction Layer and Token Position Choices
We follow Arditi et al. (2024) to select the best-performing layer and token positions for extracting the refusal direction $\mathbf{r}$. The choices are reported in Table 15.
@@ D.5 Refusal Direction Intervention with Other Model Families
Table 17: Refusal scores after ablating the instruct refusal direction from instruct (ins-ins) and abalting the base refusal direction from instruct (base-ins). The best intervention results are shown in bold. The base to instruct intervention is not very successful.
@@ Appendix E Supplementary Details and Experiments for Confidence
Table 18 shows the stats of entropy neurons across models. We observe a high overlap of entropy neurons between base and post models. To further investigate the overlapping entropy neurons, we calculate the ratio of $|\frac{\text{weight norm}}{log(\text{LogitVar})}|$ of the overlapping entropy neurons. We calculate the difference of this ratio between base and post models, and this result is also shown in Table 18. As a reference, the average ratio of all the entropy neurons is 0.0880, while the average difference of this ratio on the overlapping entropy neurons between base and post is generally less than 1% of it. It confirms that the entropy neurons are not only overlapping, but the overlapping entropy neurons are also very similar.
Model pair Overlapping neuron count (out of 10) Average ratio difference
@@
Table 18: Entropy neuron results. “Overlapping neuron count” shows the number of overlapping entropy neurons between base and post models. “Average ratio difference” shows the average difference of $|\frac{\text{weight norm}}{log(\text{LogitVar})}|$ of the overlapping entropy neurons between base and post models. As a reference, the average $|\frac{\text{weight norm}}{log(\text{LogitVar})}|$ is 0.0880 for all entropy neurons, which is much larger than the difference. base models and post models have very similar entropy neurons.

### v3

@@ 7 Confidence
Confidence of LLMs is represented by the probability associated with the decoded token. Post-trained models are known to have different confidence compared to base models (openai2024gpt4technicalreport), which is also revealed in their drastically different outputs to the same prompts. Understanding and calibrating model confidence is an important research direction. Recently, entropy neurons have been shown to be a hidden mechanism of modulating confidence that is persistent across models (gurnee2024universalneuronsgpt2language; stolfo2024confidenceregulationneuronslanguage). They have relatively high weight norms and a low composition with the model’s unembedding matrix, so they influence the model’s output probability without affecting the probability ranking much, working similarly to the temperature parameter. We study whether the difference in confidence between base and post models is caused by the difference in entropy neurons.
Entropy neuron identification Entropy neurons are identified by checking the weight norm and the variance of logit attribution. First, we compute the logit attribution for each neuron in the final MLP layer by projecting its output weights $\mathbf{w}_{\text{out}}$ onto the vocabulary space through the unembedding matrix $\mathbf{W_{U}}$. This projection approximates the neuron’s direct effect on the final prediction logits. We then calculate the variance of the normalized projection:
@@
where $\|\cdot\|_{\text{dim}=1}$ denotes a row-wise norm. A low LogitVar value indicates a relatively balanced contribution across all vocabulary tokens rather than promoting specific tokens. Entropy neurons typically have both a low LogitVar and a large weight norm to ensure they are influential. Our identification process first selects the top 25% of neurons with the largest weight norms, and from this subset, we identify the 10 neurons with the lowest LogitVar values as entropy neurons from the final MLP layer.
In our analysis comparing base and post models, we found substantial overlap in identified entropy neurons, with highly similar weight norm to LogitVar ratios. We show the detailed results in Appendix E. These findings suggest that the confidence regulation mechanism of entropy neurons remains largely unchanged during post-training, indicating that the observed confidence difference between base and post models likely stem from more subtle mechanistic changes, which require more sophisticated interpretability tools beyond entropy neurons to fully understand.
@@ 8 Discussion and Conclusion
To achieve effective post-training, it is important to understand how it shapes LLMs internally. In this paper, we analyze its effect on LLM’s internal mechanisms from four representative perspectives. We discover that post-training does not alter knowledge-storage locations and truthfulness directions significantly, and it adapts original knowledge representations while developing some new ones. In contrast, post-training changes the refusal direction, so refusal steering cannot be transferred forward. We also find that the confidence difference brought by post-training cannot be attributed to entropy neurons, which requires further investigation.
Our findings could also benefit many real-world applications. As we have shown, general abilities such as factual knowledge and the internal belief of truthfulness are mostly developed during pre-training and remain unchanged in post-training, and the forward transfer remains valid. Therefore, for fixing model errors or updating knowledge, this finding allows us to conveniently and effectively transfer knowledge edits or truthfulness probes developed on a base model to its post model. On the other hand, for internal mechanisms corresponding to abilities developed during post-training, such as refusing harmful instructions, a promising future application is to transfer the newly acquired capabilities from the post model to the base model to induce such ability without training.
Looking ahead, although we concentrated on four key perspectives, future work could extend our framework to more complex capabilities, such as reasoning and instruction-following. These areas present significant methodological challenges for existing interpretability tools. We find that properly defining the instruction-following ability is tricky, and a suitable technique to interpret this ability and verify it on base is non-trivial. Moreover, future work could leverage our analysis to improve the post-training effectiveness and efficiency.
@@ Appendix A Details on Datasets
The Tulu_extracted dataset is an in-distribution dataset for the Llama-3.1-8B sft and Mistral-7B-v0.3 sft models. In order to construct it, we use GPT-4o to extract 100 factual knowledge statements from the Tulu-SFT dataset that was used to fine-tune the sft models (Tulu). Then we use GPT-4o to generate a false statement for each true factual statement by changing the subject, object, or subject-object relation.
@@ B.2 Few-shot prompting
For each dataset, we select 2 true examples and 2 false examples to conduct four-shot prompting. We randomly select them from the dataset once, and then we fix them. The selected examples are shown in Table 7. The input is constructed using the template: “[four examples] [final statement] This statement is:”. To eliminate the influence of example order, we randomly perturb the four examples for every (true, false) statement pairs, so different pairs might have different example orders, but the true and false statements in a pair have the same example order. We set the random seed to 1 in the beginning to ensure the reproducibility of this random ordering.
@@ B.4 Supplementary Quantitative Results
Our main experiments follow the setting of marks2024geometrytruthemergentlinear. We ask the LLM to classify the truthfulness of a statement. This setup differs from the traditional causal tracing setup (meng2022locating), which uses LLM to output the object corresponding to a given subject. We choose this setting because of the following considerations. First, this setting (e.g., ”The city of Toronto is in Canada. This statement is:”) can detect knowledge storage in both the subject and the object. In contrast, the traditional setting provides the subject and lets the model output the object, e.g., ”The city of Toronto is in”. It can only detect knowledge storage in the subject. Second, our setting can test a wider range of factual knowledge. The traditional setting evaluates the patching’s influence by examining the output logit of the correct object, so it must have a fixed correct answer, such as the country of a city. But in many datasets, such as larger_than, statements like ”86 is larger than 57” don’t have a fixed correct answer. Any number less than the subject is correct here.
@@ C.3 Probe Transfer Accuracy on Mistral Family
Table 12: Probe transfer accuracy ($\uparrow$) of Mistral-7B-v0.3 base and instruct tested on 6 truthfulness datasets. For each row, we use the datasets from the other 5 rows for training. $p_{model_{1}}\rightarrow h_{model_{2}}$ means using the probe trained on $model_{1}$ to classify statement representations in $model_{2}$. Accuracy of transferred probes show little difference ($\Delta$) compared to the same-model probes.
@@ C.4 Probe Intervention Coefficient Choice
To assess the robustness of our findings to the choice of scaling factor, we extended our experiments beyond the default scalar setting ($\lambda=\pm 1$) used in marks2024geometrytruthemergentlinear. Prior work has shown that scaling can impact intervention effectiveness (li2024inferencetimeinterventionelicitingtruthful), motivating a broader evaluation.
@@ C.5 Probe Intervention on Mistral Family
Table 14: Intervention effect ($\uparrow$) of intervention on Mistral-7B-v0.3 base and instruct tested on 6 truthful datasets. For each row, we use the datasets from the other 5 rows for training. $t_{model_{1}}\mapsto h_{model_{2}}$ means using the truthfulness direction in $model_{1}$ to intervene $model_{2}$. Transfer truthful interventions show small differences ($\Delta$).
@@ D.1 Refusal Scores
Following arditi2024refusallanguagemodelsmediated; BaseLLMsRefuseToo We use refusal keywords to determine if a model refuses an input. The keywords are listed in Figure 5. Refusal score on a dataset is calculated as the percentage of outputs of the model that contains these keywords at the beginning.
@@ D.3 Refusal Direction Layer and Token Position Choices
We follow arditi2024refusallanguagemodelsmediated to select the best-performing layer and token positions for extracting the refusal direction $\mathbf{r}$. The choices are reported in Table 15.
@@ D.5 Refusal Direction Intervention with Other Model Families
Table 17: Intervention Refusal scores after ablating the instruct refusal direction from instruct (ins-ins) and abalting the base refusal direction from instruct (base-ins). The best intervention results are shown in bold. The base to instruct intervention is not very successful.
@@ Appendix E Supplementary Details and Experiments for Confidence
Table 18 shows the stats of entropy neurons across models. We observe a high overlap of entropy neurons between base and post models. To further investigate the overlapping entropy neurons, we calculate the ratio $\frac{\|\mathbf{w}_{\text{out}}\|}{log(\text{LogitVar})}$ of overlapping entropy neurons to quantitatively represent how much each neuron is qualified as an entropy neuron. We then compute the absolute difference between these ratios for entropy neurons in the base and post models with the result shown in Table 18. As a reference, the average ratio of all the entropy neurons is -0.0880, while the average absolute difference of the ratio on the overlapping entropy neurons between base and post is generally less than 1% of it. It confirms that the entropy neurons are not only overlapping, but the overlapping entropy neurons are also very similar.
Model pair Overlapping neuron count (out of 10) Avg abs ratio difference
@@
Table 18: base models and post models have very similar entropy neurons. “Overlapping neuron count” shows the number of overlapping entropy neurons between base and post models. “Avg abs ratio difference” shows the average absolute difference of $\frac{\|\mathbf{w}_{\text{out}}\|}{log(\text{LogitVar})}$ of the overlapping entropy neurons between base and post models. As a reference, the average ratio is -0.0880 for all entropy neurons.